there are three numbers and the greatest is twice the least and the middle is three more than the least and the numbers total 75. what are the numbers
Answer by DrBeeee(684)

Let the three numbers be a, b, c from least to largest.
We have a+b+c = 75
c=2a
b= a+3
By substitution we have a + (a + 3) + 2a = 75 or
4a = 75 - 3
4a = 72
a = 18
Then
b = 3+18
b = 21
and
c = 2*18
c = 36
Check. Does a+b+c=75?
Is 18+21+36 = 75?
Is 39+36 = 75?
Is 75 = 75? Yes
Your answer is 18, 21, 36